# Learning Reward Machines from Partially Observed Optimal Policies

### v1

@@ IV Examples
For our third example, we implement a modified block world problem [31, 32]. There are three blocks colored green, yellow and red, as well as 3 piles. Each stacking configuration of the blocks is a state of the MDP, and the action space consists of selecting a pile to pick from and a pile to place onto. We can only grab the top block of any stack. Action outcomes are assumed deterministic. The goal is to stack the blocks in the ordered configurations $\mathrm{a,b,c}$ shown in Figure 3. All other states have the label $\mathrm{I}$, denoting intermediate states. The corresponding reward machine is shown in Figure 4. If the robot stacks the blocks in the order $\mathrm{a}\to\mathrm{b}\to\mathrm{c}$, it gets a reward of $1$. With a depth $9$ prefix tree policy and $u_{\mathrm{max}}=3$, our algorithm recovers $2$ consistent reward machines, which are the true reward machine up-to-renaming.
@@
Figure 3: Block World Problem. The left-most stacking configuration has label $a$, where all blocks are stacked on the first pile with green being under yellow and yellow being under red. Similarly, the middle configuration has label $b$ and the right-most configuration has label $c$.
Figure 4: Stacking Reward Machine.
@@
For our last example, we introduce a “bad” state, shown in Figure 5(a). The true reward machine is shown in Figure 5(b). The robot’s task is to stack the blocks in the order $\mathrm{a}\to\mathrm{b}$ without going through $\mathrm{d}$. If it does so, it reaches $u_{2}$ and gets a reward of $1$ forever. If during execution it passes through $\mathrm{d}$, it will get a smaller (yet more immediate) reward of $0.2$, but it will get stuck at $u_{3}$ with 00 reward forever. We note that the product policy is uniformly random in both $u_{2}$ and $u_{3}$. This means that traces such as $\mathrm{a,I,b}$ and $\mathrm{a,I,d,b}$ look identical from a policy perspective, as both reach nodes with uniformly random policies, while the first being more desirable than the second. With a depth $8$ policy and $u_{\mathrm{max}}=4$, our algorithm recovers more than $1000$ consistent reward machine models, reflecting the ambiguity in this specification. However, when setting $u_{\mathrm{max}}=3$, we recover the reward machine shown in Figure 6 up-to-renaming, that is, we find a smaller reward machine consistent for this task.
@@
Figure 5: (a): the block stacking configuration with label $d$ that we want our robot to avoid. (b): the ground truth reward machine capturing that behavior.
Figure 6: Minimal consistent reward machine with the task of example 4.
V Conclusion
@@
In this work, we present a framework for learning reward machines from partially-observed optimal policies, where neither the rewards nor the reward machine states are available to the learner. Our method, instead, uses a prefix tree policy of sufficient depth in constructing a SAT problem instance from which a reward machine equivalent to the true underlying one can be obtained. Several examples demonstrate the success of our approach. In the future, we plan to extend the approach to the finite-sample settings, where the prefix tree policy is constructed from a finite demonstration set.
@@ Appendix A Proofs
A-A Proof of Proposition 1
Proof:
@@ A-A Proof of Proposition 1
Let $u\triangleq\delta_{\textbf{u}}^{*,\mathrm{learned}}(u_{I},\sigma^{\prime})$. The left-hand-side of Eq. (11) can be written as:
$\pi_{\mathrm{PTP}}^{\mathrm{learned}}(a|s,\sigma^{\prime})=\pi_{\mathrm{Prod}}^{\mathrm{learned}}(a|s,u)$
$=\pi_{\mathrm{PTP}}^{\mathrm{true}}(a|s,\sigma),\text{ for some }\sigma\in\mathrm{Path}(u)\cap\mathrm{Pref}(\mathcal{L}(\mathcal{M}_{L})),$
[…]
where the second equality is due to Eq. (10). We get:
@@
More precisely, $\delta_{\mathbf{u}}(u_{I},\sigma)=\delta_{\mathbf{u}}(u_{I},\sigma^{\prime})$. Due to the contrapositive of Lemma 1, we have a contradiction. Similarly, due to the full specification condition of Section III-A2 and Eq. (10), the support of the two prefix policies will be the same by construction. ∎
A-B Proof of Proposition 2
@@
For the $\implies$ direction, take a satisfying assignment $\{B_{k}\}_{k=1}^{|\mathrm{AP}|}$ for $\texttt{SAT}_{{u_{\mathrm{max}}},l}$ and assume by contradiction that $\{B_{k}\}_{k=1}^{|\mathrm{AP}|}$ is not satisfying for $\texttt{SAT}_{{u_{\mathrm{max}}},j}$ with $j>l$. Consider the reward machine model $\mathcal{G}^{\mathrm{learned}}$, with the transition function $\delta_{\textbf{u}}^{\mathrm{learned}}$, corresponding to $\{B_{k}\}_{k=1}^{|\mathrm{AP}|}$. Since $\{B_{k}\}_{k=1}^{|\mathrm{AP}|}$ is unsatisfying for $\texttt{SAT}_{{u_{\mathrm{max}}},j}$, then there exists a negative example $(\sigma,\sigma^{\prime})\in\mathcal{E}^{-}_{j}\setminus\mathcal{E}^{-}_{l}$ such that
@@
Hence, by Eq. (12), we have that $u^{\mathrm{learned}}=u^{\mathrm{learned},\prime}$. Let $\mathcal{G}^{\mathrm{sync}}$ be the synchronized reward machine model between $\mathcal{G}^{\mathrm{learned}}$ and $\mathcal{G}^{\mathrm{true}}$ according to Definition 3. We observe the following:
$(u^{\mathrm{true}},u^{\mathrm{learned}})$ $=\delta_{\mathbf{u}}^{\mathrm{sync},*}(u_{I}^{\mathrm{sync}},\sigma),$
$(u^{\mathrm{true}},u^{\textsf{learned}},s^{*})\in\mathcal{G}^{\textsf{sync}}\times\mathcal{M}_{L},$
$(u^{\textsf{true},\prime},u^{\textsf{learned},\prime})$ $=\delta_{\mathbf{u}}^{\textsf{sync},*}(u_{I}^{\textsf{sync}},\sigma^{\prime}),$
$(u^{\textsf{true},\prime},u^{\textsf{learned},\prime},s^{*})\in\mathcal{G}^{\textsf{sync}}\times\mathcal{M}_{L}.$
@@
Note that $s^{*}$ is still the reached MDP state in both synchronized product nodes above. By the definition of $\delta_{\mathbf{u}}^{\textsf{sync},*}$, we can decompose Eq. (A-B) into:
@@
Since $\pi_{\mathrm{PTP}}(a^{*}|s^{*},\sigma)\neq\pi_{\mathrm{PTP}}(a^{*}|s^{*},\sigma^{\prime})$ due to our contradiction assumption, we conclude from Eq. (A-B) that $\pi_{\mathrm{PTP}}(a^{*}|s^{*},\bar{\sigma})\neq\pi_{\mathrm{PTP}}(a^{*}|s^{*},\bar{\sigma}^{\prime})$. However, $u^{\mathrm{learned}}=u^{\mathrm{learned},\prime}$ combined with Eq. (A-B) implies that $\delta_{\textbf{u}}^{\mathrm{learned}}(u_{I},\bar{\sigma})=\delta_{\textbf{u}}^{\mathrm{learned}}(u_{I},\bar{\sigma}^{\prime})$, contradicting that $\{B_{k}\}_{k=1}^{|\mathrm{AP}|}$ is a SAT assignment for the depth $l$. This concludes the proof. ∎

### v2

@@ 5.2 Tabular BlockWorld MDPs
The setup for this experiment is a modified block world problem (Khodeir et al., 2023; Wolfe and Barto, 2006). There are three blocks colored green, yellow and red, as well as 3 piles. Each stacking configuration of the blocks is a state of the MDP, and the action space consists of selecting a pile to pick from and a pile to place onto. We can only grab the top block of any stack. Action outcomes are assumed deterministic. The goal in the first task (stack) is to stack the blocks in the ordered stacking configurations $\mathbf{st_{1}},\mathbf{st_{2}},\mathbf{st_{3}}$, shown in Figure 2(a). All other states have the label $\mathbf{i}$, denoting intermediate states. The corresponding reward machine is shown in Figure 2(b). If the robot stacks the blocks in the order $\mathbf{st_{1}}\to\mathbf{st_{2}}\to\mathbf{st_{3}}$, it gets a reward of $1.0$. With a depth-$9$ prefix tree policy and $u_{\mathrm{max}}=3$, our algorithm recovers $\mathbf{2}$ consistent reward machines, which are the true reward machine up-to-renaming. The findings are summarized in Table 2. Results obtained from a finite set of demonstrations are reported in Appendix B.1.1.
@@
Figure 2: (a): Block World MDP. The left-most stacking configuration has label $\mathbf{st_{1}}$, where all blocks are stacked on the first pile with green being under yellow and yellow being under red. Similarly, the middle configuration has label $\mathbf{st_{2}}$ and the right-most configuration has label $\mathbf{st_{3}}$. (b): Stacking Reward Machine.
@@
For our second task (stack-avoid), we introduce a “bad” state, shown in Figure 3(a). The true reward machine is shown in Figure 3(b). The robot’s task is to stack the blocks in the order $\mathbf{st_{1}}\to\mathbf{st_{2}}$ without going through ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}}$. If it does so, it reaches $u_{2}$ and gets a reward of $1$ forever. If during execution it passes through ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}}$, it will get a smaller (yet more immediate) reward of $0.2$, but it will get stuck at $u_{3}$ with 00 reward forever. We note that the product policy is uniformly random in both $u_{2}$ and $u_{3}$. This means that proposition traces such as $\mathbf{st_{1},i,st_{2}}$ and $\mathbf{st_{1},i,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}},st_{2}}$ look identical from a policy perspective, as both reach nodes with uniformly random policies, while the first being more desirable than the second. By setting $u_{\mathrm{max}}=3$, a depth-$8$ policy was sufficient to recover the reward machine shown in Figure 3(c) up-to-renaming; that is, we find a smaller reward machine consistent for this task. Notably, setting $u_{\mathrm{max}}=4$ with the same depth policy yields more than a 1000 solutions, indicating that uniquely recovering the same sized reward machine -as the ground truth machine- requires longer depth policies. Results obtained from a finite set of demonstrations are reported in Appendix B.1.2.
@@
Figure 3: (a) The block stacking configuration with label ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}}$ that we want our robot to avoid. (b) The ground truth reward machine. (c) Smaller consistent reward machine with the task stack-avoid.
Task Depth $|\mathcal{S}|$ $|\mathcal{A}|$ $|\mathcal{E}^{-}|$ SAT time (sec)
stack 9 60 9 31656 0.558
stack-avoid 8 60 9 24763 0.483
Table 2: Solution statistics for the tabular BlockWorld MDP.
5.3 2D-Continous Reacher Robotic Arm
Our setup for this experiment is a modified Reacher-v5 environment (Gymnasium (Towers et al., 2024); MuJoCo (Todorov et al., 2012)), where a 2D robotic arm must reach targets randomly placed within the arena. The agent’s state is a 10-D continuous vector consisting of the end-effector’s position and velocity, the target’s coordinates, and their mutual distance. Actions are continuous torques in $[-1,1]^{2}$ applied at the elbow and shoulder joints.
Inspired by (Araki et al., 2021), we fix three colored targets — blue ($\mathbf{b}$), red ($\mathbf{r}$), and yellow ($\mathbf{y}$) — in distinct quadrants of the arena (see Figure 4). The goal is to reach them in the order $\mathbf{b}\to\mathbf{y}\to\mathbf{r}$. To train an expert policy, we discretize each torque dimension into five levels $\{-1.0,-0.5,0.0,0.5,1.0\}$ (yielding 25 total actions). We then employ Proximal Policy Optimization (PPO) (Schulman et al., 2017), as implemented in Stable-Baselines3 (Raffin et al., 2021), to maximize a reward given by the negative Euclidean distance between the end-effector and the active target.
When generating trajectories, we switch the desired target immediately upon reaching the previous target, thereby emulating a reward machine without introducing the machine during training or simulation. Appendix B.2 provides full discretization and training details. From this procedure we collect $1M$ trajectories.
Table 3 reports, for various probability thresholds $\alpha$ (defined in Section 4.2), the size of the negative-example set $\mathcal{E}^{-}$ and its False Positive Rate (FPR) [^1] FPR is simply the number of false positives divided by $|\mathcal{E}^{-}|$.. As expected from Section 4, solving our SAT problem with all the negative examples yields $\mathbf{0}$ solutions. Instead, we employ our weighted MAX-SAT variant and recover two viable solutions (Figures 4(b) and 4(c)) up-to-renaming, which represent the best achievable given our dataset. These are the same solutions that we recover if instead we supervised with the ground-truth reward machine to remove the false positives, emphasizing the robustness of this approach.
Figure 4: Reacher experiment. (a): 2-link robotic arm with the three colored targets. (b) First recovered reward machine model. (c) Second recovered reward machine model.
$|\mathfrak{D}|$ $|\mathcal{S}|$ $|\mathcal{A}|$ $|\tau|$ $|\mathcal{E}^{-}|$ $\alpha$ FPR # Weighted MAX-SAT solutions
1M 17.4M 25 160 1472 0.001 1.90% 4
1M 17.4M 25 160 1193 0.0001 1.67% 4
1M 17.4M 25 160 882 0.00001 1.36% 4
Table 3: Summary of the experiments for the 2D-continuous reacher arm.
5.4 Real-world Mice Navigation
We also applied our learning framework to the trajectories of real mice navigating (in the dark) in a 127-node labyrinth maze (Rosenberg et al., 2021) shown in Figure 5. Each node, labeled with a number in Figure 5(a), represents a state in the MDP. The mouse can select from 4 actions: $\{\mathrm{stay},\mathrm{right},\mathrm{left},\mathrm{reverse}\}$. State $116$ (middle-right) contains a water resource and is labeled $\mathbf{w}$, state 00 (center) is the home state labeled $\mathbf{h}$, and all other states are labeled $\mathbf{i}$. Two cohorts of 10 mice moved freely in the maze for 7 hours, with one cohort being water-restricted and the other was not. A sample water-restricted mouse trajectory is shown in Figure 5(b). Difference in water restriction condition resulted in different animal behavior between these two cohorts. For the purposes of our study, we only considered the water-restricted mice. We used the same dataset of trajectories from (Ashwood et al., 2022), which is comprised of 200 mouse trajectories, given as state-action pairs of length $22$ each. We set our probability threshold at $\alpha=0.001$. With $u_{\mathrm{max}}=2$, our algorithm learns the unique reward machine shown in Figure 5(c). This is consistent with the seen behaviors of water-restricted mice which first try to reach the water port and hydrate, after which their behaviors switch to exploring the maze or going back the home state (Zhu et al., 2024; Ashwood et al., 2022). More details can be found in Appendix B.3.
Figure 5: Labyrinth experiment. (a): Maze structure and state space definition. (b): Trajectory of a single mouse. (c): Recovered reward machine. (a) and (b) are reprinted from (Rosenberg et al., 2021).
6 Conclusion
@@
In this work, we present a framework for learning reward machines from partially observed policies, where neither the rewards nor the reward machine states are available to the learner. Instead, our method constructs a SAT problem instance from a sufficiently deep prefix tree policy, from which a reward machine equivalent to the true underlying one can be recovered. The approach is generalized to learn the reward machines directly from demonstrations from an optimal expert, where robustness to inaccuracies in policy estimates is handled using finite-sample confidence bounds and MAX-SAT is used to eliminate incorrect negative examples. A diverse set of experiments illustrate the effectiveness of our approach.
@@ Appendix A Proofs
A.1 Proof of Proposition 1
Proof.
@@ A.1 Proof of Proposition 1
Let $u\triangleq\delta_{\textbf{u}}^{*,\mathrm{learned}}(u_{I},\sigma^{\prime})$. The left-hand-side of (13) can be written as:
$\pi_{\mathrm{PTP}}^{\mathrm{learned}}(a|s,\sigma^{\prime})=\pi_{\mathrm{Prod}}^{\mathrm{learned}}(a|s,u)=\pi_{\mathrm{PTP}}^{\mathrm{true}}(a|s,\sigma),\text{ for some }\sigma\in\mathrm{Path}(u)\cap\mathrm{Pref}(\mathcal{L}(\mathcal{M}_{L})),\text{ with }|\sigma|\leq l,$
where the second equality is due to (10). We get:
@@
More precisely, $\delta_{\mathbf{u}}(u_{I},\sigma)=\delta_{\mathbf{u}}(u_{I},\sigma^{\prime})$. Due to the contrapositive of Lemma 1, we have a contradiction. Similarly, due to the full specification condition of Section 3.1.2 and (10), the support of the two prefix policies will be the same by construction. ∎
A.2 Proof of Proposition 2
@@
For the $\implies$ direction, take a satisfying assignment $\{B_{k}\}_{k=1}^{|\mathrm{AP}|}$ for $\texttt{SAT}_{{u_{\mathrm{max}}},l}$ and assume by contradiction that $\{B_{k}\}_{k=1}^{|\mathrm{AP}|}$ is not satisfying for $\texttt{SAT}_{{u_{\mathrm{max}}},j}$ with $j>l$. Consider the reward machine model $\mathcal{G}^{\mathrm{learned}}$, with the transition function $\delta_{\textbf{u}}^{\mathrm{learned}}$, corresponding to $\{B_{k}\}_{k=1}^{|\mathrm{AP}|}$. Since $\{B_{k}\}_{k=1}^{|\mathrm{AP}|}$ is unsatisfying for $\texttt{SAT}_{{u_{\mathrm{max}}},j}$, then there exists a negative example $\{\sigma,\sigma^{\prime}\}\in\mathcal{E}^{-}_{j}\setminus\mathcal{E}^{-}_{l}$ such that
@@
Hence, by (14), we have that $u^{\mathrm{learned}}=u^{\mathrm{learned},\prime}$. Let $\mathcal{G}^{\mathrm{sync}}$ be the synchronized reward machine model between $\mathcal{G}^{\mathrm{learned}}$ and $\mathcal{G}^{\mathrm{true}}$ according to Definition 3. We observe the following:
$(u^{\mathrm{true}},u^{\mathrm{learned}})$ $=\delta_{\mathbf{u}}^{\mathrm{sync},*}(u_{I}^{\mathrm{sync}},\sigma),\quad(u^{\mathrm{true}},u^{\textsf{learned}},s^{*})\in\mathcal{G}^{\textsf{sync}}\times\mathcal{M}_{L},$
$(u^{\textsf{true},\prime},u^{\textsf{learned},\prime})$ $=\delta_{\mathbf{u}}^{\textsf{sync},*}(u_{I}^{\textsf{sync}},\sigma^{\prime}),\quad(u^{\textsf{true},\prime},u^{\textsf{learned},\prime},s^{*})\in\mathcal{G}^{\textsf{sync}}\times\mathcal{M}_{L}.$
@@
Note that $s^{*}$ is still the reached MDP state in both synchronized product nodes above. By the definition of $\delta_{\mathbf{u}}^{\textsf{sync},*}$, we can decompose (A.2) into:
@@
Since $\pi_{\mathrm{PTP}}(a^{*}|s^{*},\sigma)\neq\pi_{\mathrm{PTP}}(a^{*}|s^{*},\sigma^{\prime})$ due to our contradiction assumption, we conclude from (A.2) that $\pi_{\mathrm{PTP}}(a^{*}|s^{*},\bar{\sigma})\neq\pi_{\mathrm{PTP}}(a^{*}|s^{*},\bar{\sigma}^{\prime})$. However, $u^{\mathrm{learned}}=u^{\mathrm{learned},\prime}$ combined with (A.2) implies that $\delta_{\textbf{u}}^{\mathrm{learned}}(u_{I},\bar{\sigma})=\delta_{\textbf{u}}^{\mathrm{learned}}(u_{I},\bar{\sigma}^{\prime})$, contradicting that $\{B_{k}\}_{k=1}^{|\mathrm{AP}|}$ is a SAT assignment for the depth $l$. This concludes the proof. ∎
@@
Appendix B Experiment Details and Additional Results
B.1 Tabular BlockWorld MDP
To generate the expert trajectory dataset, we first compute the product policy given the BlockWorld MDP dynamics and the corresponding ground-truth reward machine. We use soft bellman iteration to find the optimal product policy. The procedure is summarized in Algorithm 2.
Input: MDP $\mathcal{M}$, ground-truth reward machine $\mathcal{R}$, labeling function $L$, tolerance $\epsilon$.
Output: Optimal product policy $\pi^{*}$
1
2Construct product MDP $\mathcal{M}_{\mathrm{Prod}}=\bigl{(}\mathcal{S}^{\prime},\;\mathcal{A}^{\prime},\;\mathcal{P}^{\prime},\;\mu_{0}^{\prime},\;\gamma^{\prime},\;r^{\prime}\bigr{)}$ as in Section 2.2.
3Initialize
$V_{0}(s,u)\leftarrow 0\quad\forall(s,u)\in\mathcal{S}^{\prime},\quad\text{error}\leftarrow\infty,\quad k\leftarrow 0$
4while $\text{error}>\epsilon$ do
5       foreach $(s,u)\in\mathcal{S}^{\prime}$ do
6             foreach $a\in\mathcal{A}^{\prime}$ do
7                   $Q_{k+1}(s,u,a)\;\leftarrow\;r^{\prime}\bigl{(}(s,u),a\bigr{)}\;+\;\gamma^{\prime}\;\sum_{(s^{\prime},u^{\prime})\in\mathcal{S}^{\prime}}\mathcal{P}^{\prime}\bigl{(}(s,u),a,(s^{\prime},u^{\prime})\bigr{)}\,V_{k}(s^{\prime},u^{\prime})$
8            $V_{k+1}(s,u)\;\leftarrow\;\log\!\displaystyle\sum_{a\in\mathcal{A}^{\prime}}\exp\!\bigl{(}Q_{k+1}(s,u,a)\bigr{)}$
9      $\text{error}\;\leftarrow\;\max_{(s,u)\in\mathcal{S}^{\prime}}\bigl{|}V_{k+1}(s,u)-V_{k}(s,u)\bigr{|}$
10       $k\;\leftarrow\;k+1$
11
12foreach $(s,u)\in\mathcal{S}^{\prime},\,a\in\mathcal{A}^{\prime}$ do
13       $\pi^{*}(a\mid s,u)\;\leftarrow\;\exp\!\bigl{(}Q_{k}(s,u,a)-V_{k}(s,u)\bigr{)}$
14
return $\pi^{*}$
Algorithm 2 Soft Bellman Iteration on the Product MDP
At each time step, we use the visited proposition prefix $\sigma$ to find the reached node $u$ on the true reward machine, using $u=\delta_{\mathrm{u}}^{*}(u_{I},\sigma)$. We then sample an action from the true product policy $\pi^{*}(a|s,u)$, where $s$ is the MDP state reached. We keep count of the sampled actions at all state-prefix pairs for constructing the prefix tree policy. The procedure is summarized in Algorithm 3.
Input: MDP $\mathcal{M}$, reward machine $\mathcal{R}$, true product policy $\pi^{*}(a\mid s,u)$, # trajectories $N$, trajectory length $H$
Output: Prefix-tree policy $\pi_{\mathrm{PTP}}^{H}$
1
2Initialize counts: $C(\sigma,s,a)\leftarrow 0$ for all prefixes $\sigma$, states $s\in\mathcal{S}$, actions $a\in\mathcal{A}$.
3for $i\leftarrow 1$ to $N$ do
4       $\sigma\leftarrow\epsilon$ (empty prefix)
5       $s\sim\mu_{0}$ (initial MDP state)
6       $u\leftarrow u_{I}$ (initial RM node)
7       for $t\leftarrow 1$ to $H$ do
8             $u\leftarrow\delta_{\mathrm{u}}(u,\sigma)$;
9             Sample $a\sim\pi^{*}(\cdot\mid s,u)$;
10             Execute $a$, observe $s^{\prime}\sim\mathcal{P}(s,a)$ and label $\ell\leftarrow L(s^{\prime})$;
11             $\sigma\leftarrow\sigma\,\|\,\ell$  (concatenation);
12             $C(\sigma,s,a)\leftarrow C(\sigma,s,a)+1$;
13             $s\leftarrow s^{\prime}$;
14
15
16
17foreach prefix $\sigma$, state $s$, action $a$ do
18       $\pi_{\mathrm{PTP}}^{H}(a\mid s,\sigma)\;=\;\frac{C(\sigma,s,a)}{\sum_{a^{\prime}\in\mathcal{A}}C(\sigma,s,a^{\prime})}\,.$
19
return $\pi_{\mathrm{PTP}}^{H}$
Algorithm 3 Constructing the Prefix-Tree Policy via Simulation
In the following section, we show the results for the stack and stack-avoid tasks.
B.1.1 (stack) task
We simulated varying size demonstration datasets and solved our SAT problem with each. In Table 4, we show the effect of the number of demonstrations in reducing the number of satisfying solutions. With a demonstration dataset of size $1M$, the ground-truth (stack) reward machine is recovered up-to-renaming. However, even a demonstration set of size $5000$ can indeed reduce the number of satisfying solutions to $4$. These $4$ solutions include the ground-truth reward machine (accounting for $2$ solutions due to renaming), and the reward machine (up-to-renaming again) shown in Figure 7(b), where $\delta_{\mathrm{u}}(u_{2},\mathbf{a})=u_{1}$, instead of $\delta_{\mathrm{u}}(u_{2},\mathbf{a})=u_{2}$ as in the ground-truth reward machine.
$|\mathfrak{D}|$ $|\tau|$ $|\mathcal{E}^{-}|$ $\alpha$ SAT solutions
1000 20 84 0.05 24
3000 20 226 0.05 12
5000 20 304 0.05 4
10000 20 456 0.05 4
0.1M 20 1965 0.05 4
1M 20 6242 0.05 2
Table 4: Size of the simulated dataset vs. the number of satisfying solutions.
Figure 7: Recovered Reward Machine Models for stack task.
B.1.2 (stack-avoid) task
In Table 5, we similarly show the effect of the size of the simulated dataset on the number of satisfying solutions for the stack-avoid task. In this task, $1M$ trajectories were not enough to extract all the necessary negative examples to reduce the number of solutions to $2$. However, a much smaller dataset of size $1000$ was enough to reduce the number of solutions to $\mathbf{4}$, which we can manually inspect. These solutions are shown in Figures 8(a) and  8(b) (up-to-renaming). The reward machine in Figure 8(b) differs from the ground-truth reward machine by assigning $\delta_{\mathrm{u}}(u_{1},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}})=u_{1}$, instead of $\delta_{\mathrm{u}}(u_{1},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}})=u_{2}$. Such an ambiguity could have been resolved by a negative example of the form $\{\sigma_{1}=\mathbf{st_{1},i,st_{1}},\sigma_{2}=\mathbf{st_{1},i,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}},i,st_{1}}\}$, which the MDP is capable of producing given the length of the simulated trajectories. However, since the policy becomes uniformly random after reaching ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}}$, simulating such negative examples might require a much larger dataset.
$|\mathfrak{D}|$ $|\tau|$ $|\mathcal{E}^{-}|$ $\alpha$ SAT solutions
100 20 24 0.05 40
200 20 37 0.05 8
500 20 77 0.05 8
1000 20 130 0.05 4
0.1M 20 1680 0.05 4
1M 20 4771 0.05 4
Table 5: Size of the simulated dataset vs. the number of satisfying solutions.
Figure 8: Recovered Reward Machine Models for stack-avoid task.
B.2 2D-Continuous Robotics Reacher Arm
B.2.1 Expert policy training using deep RL
To train the policy, we used the standard Reacher-v5 environment from Gymnasium, augmented with custom wrappers to discretize the action space and introduce randomized initial conditions. We also increased the episode horizon from the default $50$ steps to $160$ steps in order to match the expected time to visit all 3 desired poses and finish the task. The DiscreteReacherActionWrapper maps the continuous two-dimensional action space to a MultiDiscrete space of 25 torque combinations by restricting each component to one of five values: {-1.0, -0.5, 0.0, 0.5, 1.0}. Additionally, the ForceRandomizedReacher wrapper introduces randomized initial joint positions and velocities at each episode reset to promote robustness and generalization. This modification was essential, as the default initial conditions in the base environment are typically limited to small perturbations around fully-extended arm configurations. Training solely from such narrow initializations led to poor task performance and limited the policy’s ability to generalize across the broader state space.
We trained a Proximal Policy Optimization (PPO) agent using the stable-baselines3 library, with all the default settings and hyperparameters. To accelerate training, we employed a vectorized environment with 50 parallel instances running on CPU. The agent was trained for 20 million time steps using a multilayer perceptron (MLP) policy.
B.2.2 Discretizing the state space and sampling the policy
The MDP state space is generated by discretizing the shoulder angle $\theta_{1}$, the elbow angle $\theta_{2}$, and the angular velocities $\dot{\theta}_{1},\dot{\theta}_{2}$. The angles are discritized uniformly in $[-\pi,\pi]$ with a bin size of $10^{\circ}\approx 0.17$ radians. The angular velocities are discretized uniformly in $[-14,14]$ with a bin size of $0.25$ rad/s. This leads to $\approx 17.4M$ states in the MDP.
When sampling an action from the optimal policy, we first identify the discrete state corresponding to the current continuous state. The action is then sampled at the center point of that discrete region. This approach standardizes the policy’s behavior across different continuous states that map to the same discrete representation, ensuring consistency within each discretized region of the state space.
B.3 Real-world Mice Navigation
For recovering the reward function, we first need to construct the product policy. However, real-world data generally do not perfectly fit the assumed mathematical models, i.e., the max entropy assumption. Since most of the states in the MDP are unvisited, and many action are not sampled at most of the states, the product policy is predominantly sparse ($\sim 92\%$ of the action probabilities are 00). We preprocess the policy by clipping its minimum value to $0.05$. This allows us to use the results from [Shehab et al., 2024] to compute the reward function, as we requires computing the $\log$ of the policy. The reward space is constrained to satisfy the feature-decomposition which the reward machine provides. However, there is generally no guarantee that we can find such a featurized reward function that can perfectly induce any product policy. With the shown reward in Figure 5(c), the norm difference between the induced policy and the true policy is $2.66$ (the norm difference with a uniformly random policy is $2.97$). By increasing $u_{\mathrm{max}}$ to $3$, one of the recovered reward machines is shown in Figure 9, which yields a slightly larger policy difference norm of $2.71$.
Figure 9: Recovered Reward machine for Mice Dataset with $u_{\mathrm{max}}=3$.